# Relativistic fluid dynamics in a ‘hydro’ frame

### v1

@@ Abstract
Naively, it appears that there are $(N_{1}^{{\rm snd}}+2)$ equations for six unknown functions. However, $R_{T}$ and $R_{u}$ depend only on $k$. Taking the limit $\omega\rightarrow 0$ in the two equations of (16) effectively yields two additional equations involving $R_{T}$, $R_{u}$, and $[F_{T}~{}G_{u}]_{(0)}$ (which is independent of $\omega$ and defined as $[F_{T}~{}G_{u}]_{(0)}\equiv\lim_{\omega\rightarrow 0}[F_{T}(\omega,k)G_{u}(\omega,k)]$).
@@
Using these two additional equations, we can uniquely determine $R_{T}$ and $R_{u}$ in terms of $[F_{T}~{}G_{u}]_{(0)}$ and other known functions of $k$, such as ${\mathfrak{w}}^{a}_{{\rm snd}}(k)$ and ${\mathfrak{w}}^{a}_{\rm sh}(k)$. Next, we substitute the solutions for $R_{T}$ and $R_{u}$ and reintroduce $\omega$ into (16). We can now uniquely solve for $F_{u}$ and $G_{T}$ in terms of the product $F_{T}G_{u}$ and $[F_{T}G_{u}]_{(0)}$. At this point, both equations in (16) will be solved for all values of $\omega$ and $k$. Next, we have $N_{1}^{{\rm snd}}$ equations from (17) to determine $F_{T}$ and $G_{u}$. For an uncharged hydrodynamic theory like those discussed here, $N_{1}^{{\rm snd}}=2$, giving us two equations for the two remaining unknowns, $F_{T}$ and $G_{u}$. However, the equations in (17) are evaluated at $\omega=\omega^{a}_{{\rm snd}}(k)$, which prevents us from uniquely determining $F_{T}$ and $G_{u}$ for all values of $\omega$. In particular, we can add two arbitrary functions of the product $\prod_{a}(\omega-\omega^{a}_{{\rm snd}})$ to any solutions for $F_{T}$ and $G_{u}$, and they will still satisfy the equations in (17). Despite the minor ambiguity in $F_{T}$ and $G_{u}$ noted above, we can nearly uniquely determine the functions in the linearized frame transformation by imposing that the spectrum in the transformed frame contains only the hydrodynamic modes [28].
@@
Field redefinition and the non-hydrodynamic mode in BDNK theory:– Let us now apply the above algorithm to eliminate the non-hydrodynamic modes in the BDNK theory [10, 11]. The stress-energy tensor for an uncharged conformal system reads
@@
with $q^{\mu}\equiv(u\cdot\partial)u^{\mu}+\left(P^{\mu\nu}\partial_{\nu}T\over T\right)$, $E\equiv{(u\cdot\partial)T\over T}+{\left(\partial\cdot u\right)}/{3}$, and $\sigma^{\mu\nu}\equiv\left({P^{\mu\alpha}P^{\nu\beta}+P^{\mu\beta}P^{\nu\alpha}\over 2}-{P^{\mu\nu}P^{\alpha\beta}\over 3}\right)\left(\partial_{\alpha}u_{\beta}\right)$. Here, $\{\chi,\theta,\lambda\}$ are constants and are functions of $T$. For above tensor, the dispersion polynomial is [^1] We have re-scaled the transport coefficients ($\theta$, $\chi$, and $\lambda$) by proper factors of $T$ so that the products of these coefficients with frequency ($\omega$) or momentum ($k$) are dimensionless.
$P_{\text{BDNK}}(\omega,k)=\left[\omega(1-i\chi~{}\omega)+i\lambda k^{2}\right]$ (19)
@@
$\begin{split}{\mathfrak{w}}_{\rm sh}=&-{i\over\theta}+i\lambda~{}k^{2}\,,\quad\omega_{\rm sh}=-i\lambda~{}k^{2}\,,\\ {\mathfrak{w}}^{(1)}_{{\rm snd}}=&-{i\over\theta}+{i\over 3}\left[4\lambda-{\theta\chi\over(\theta-\chi)}\right]k^{2}\,,\\ {\mathfrak{w}}^{(2)}_{{\rm snd}}=&-{i\over\chi}+{i\over 3}\left(\theta\chi\over\theta-\chi\right)k^{2}\,,\\ \omega^{(1)}_{{\rm snd}}=&{k\over\sqrt{3}}-\left(2i\lambda\over 3\right)k^{2}-\left(2\lambda^{2}\over 3\sqrt{3}\right)k^{3}\,,\\ \omega^{(1)}_{{\rm snd}}=&-{k\over\sqrt{3}}-\left(2i\lambda\over 3\right)k^{2}+\left(2\lambda^{2}\over 3\sqrt{3}\right)k^{3}\,,\\ {\cal R}(\omega,k)=&\,k\left[i+(\theta+\chi)\omega\over k^{2}\theta+3\omega(i+\chi\omega)\right].\end{split}$ (20)
@@
$\begin{split}R_{u}(k)=\frac{1}{k^{2}}\left({\mathfrak{w}}_{\rm sh}+{i\over\theta}\over{\mathfrak{w}}_{\rm sh}\right).\end{split}$ (21)
@@
$\begin{split}F_{u}(\omega,k)=-{1/\theta\over{\mathfrak{w}}_{\rm sh}(\omega-{\mathfrak{w}}_{\rm sh})}.\end{split}$ (22)
@@
$R_{T}(k)=i\theta\,{\mathfrak{w}}_{\rm sh}\left[F_{T}\,G_{u}\right]_{(0)}+\frac{1}{k^{2}}\left[1+{i\over\chi}\left({\mathfrak{w}}_{\rm sh}\over{\mathfrak{w}}_{{\rm snd}}^{(1)}{\mathfrak{w}}_{{\rm snd}}^{(2)}\right)\right],$ (23)
with $\left[F_{T}~{}G_{u}\right]_{(0)}\equiv\lim_{\omega\rightarrow 0}\left[F_{T}(\omega,k)G_{u}(\omega,k)\right]$. Finally, putting (23) in the same equation with $\omega$ finite, we solve for $G_{T}(\omega,k)$
@@
Note that (25) must be fulfilled only at the frequencies of the hydrodynamic sound modes. Replace all explicit $\omega$ dependence with $\omega^{(1)}_{{\rm snd}}(k)$ and $\omega^{(2)}_{{\rm snd}}(k)$. Solving these equations does not determine the unknown functions $F_{T}$ and $G_{u}$ for all $\omega$ and $k$. However, our aim is to demonstrate the existence of at least one frame transformation that eliminates the non-hydrodynamic modes. By assuming $F_{T}$ and $G_{u}$ do not depend on $\omega$, equation (25) uniquely solves these functions, providing a suitable frame transformation. With this assumption, the solutions for $R_{T}$ and $G_{T}$ simplify
@@
Even after assuming $F_{T}$ and $G_{u}$ are independent of $\omega$, eq. (25) remains a set of coupled nonlinear algebraic equations, with nonlinearity due to $R_{T}$ depending on the product of $F_{T}$ and $G_{u}$. While solving is straightforward, the exact expressions for $F_{T}$ and $G_{u}$ are cumbersome compared to other functions $F_{u}$, $R_{u}$, $R_{T}$, and $G_{T}$. The intermediate details are presented in [28]. The final leading solutions, up to $\mathcal{O}(k^{2})$, are as follows:
@@
At this point, we note that although we provided formal exact solutions for $F_{u}$, $R_{u}$, $R_{T}$, and $G_{T}$, these should be understood as expansions in non-negative powers of $\omega$ and $k$. Polynomials of $k$ and $\omega$ in the denominator imply an infinite power series with a convergence radius computed by the zeros of these polynomials. Our study relies on the derivative expansion, and replacing $\{\omega,k^{2}\}$ with derivatives doesn’t make sense if $\omega$ and $k$ appear in the denominator. All functions $F_{u}$, $F_{T}$, $G_{u}$, $G_{T}$, $R_{T}$, and $R_{u}$ remain finite as $\omega\to 0$ and/or $k\to 0$. This ensures no negative powers of $(u\cdot\partial)$ or $(P_{\mu}^{\nu}\partial_{\nu})$ arise after expansion and replacement. Furthermore, the final frame transformation solution must depend only on even powers of $k$ to apply the replacement without using non-analytic expressions like the square root of a derivative. Since we’re concerned with the spectrum of linearized perturbations, the order of derivatives $(u\cdot\partial)$ and $P^{\mu}_{\nu}\partial_{\mu}$ is irrelevant. The frame transformation incorporates terms of all orders in the derivative expansion, thus, the final transformed stress-energy tensor also have terms of all orders in the derivative expansion.

### v2

@@ Abstract
Let’s examine whether we can obtain a solution for this set of coupled nonlinear equations ((16) and (17)) and to what degree the solution would be unique. Naively, it appears that there are $(N_{1}^{{\rm snd}}+2)$ equations for six unknown functions. However, $R_{T}$ and $R_{u}$ depend only on $k$. Taking the limit $\omega\rightarrow 0$ in the two equations of (16) effectively yields two additional equations involving $R_{T}$, $R_{u}$, and $[F_{T}\,G_{u}]_{(0)}$ (defined as $[F_{T}\,G_{u}]_{(0)}\equiv\lim_{\omega\rightarrow 0}[F_{T}(\omega,k)G_{u}(\omega,k)]$).
@@
Using these additional equations, we can uniquely determine $R_{T}$ and $R_{u}$ in terms of $[F_{T}~{}G_{u}]_{(0)}$ and other known functions of $k$, such as ${\mathfrak{w}}^{a}_{{\rm snd}}(k)$ and ${\mathfrak{w}}^{a}_{\rm sh}(k)$. Next, we substitute the solutions for $R_{T}$ and $R_{u}$ and reintroduce $\omega$ into (16). We can now uniquely solve for $F_{u}$ and $G_{T}$ in terms of the product $F_{T}G_{u}$ and $[F_{T}G_{u}]_{(0)}$. At this point, both equations in (16) will be solved for all values of $\omega$ and $k$. Next, we have $N_{1}^{{\rm snd}}$ equations from (17) to determine $F_{T}$ and $G_{u}$. For an uncharged hydrodynamic theory like those discussed here, $N_{1}^{{\rm snd}}=2$, giving us two equations for the two remaining unknowns, $F_{T}$ and $G_{u}$. However, the equations in (17) are evaluated at $\omega=\omega^{a}_{{\rm snd}}(k)$, which prevents us from uniquely determining $F_{T}$ and $G_{u}$ for all values of $\omega$. In particular, we can add two arbitrary functions of the product $\prod_{a}(\omega-\omega^{a}_{{\rm snd}})$ to any solutions for $F_{T}$ and $G_{u}$, and they will still satisfy (17). Despite the minor ambiguity in $F_{T}$ and $G_{u}$ noted above, we can nearly uniquely determine the functions in the linearized frame transformation by imposing that the spectrum in the transformed frame contains only the hydrodynamic modes[29].
@@
Field redefinition and non-hydrodynamic mode in BDNK:– Let us now apply the above algorithm to eliminate the non-hydrodynamic modes in the BDNK theory[9, 10]. The stress-energy tensor for an uncharged conformal system reads
@@
with $q^{\mu}\equiv(u\cdot\partial)u^{\mu}+\left(P^{\mu\nu}\partial_{\nu}T\over T\right)$, $E\equiv{(u\cdot\partial)T\over T}+{\left(\partial\cdot u\right)}/{3}$, and $\sigma^{\mu\nu}\equiv\left({P^{\mu\alpha}P^{\nu\beta}+P^{\mu\beta}P^{\nu\alpha}\over 2}-{P^{\mu\nu}P^{\alpha\beta}\over 3}\right)\left(\partial_{\alpha}u_{\beta}\right)$. Here, $\{\chi(T),\theta(T),\lambda(T)\}$ are constants. The dispersion polynomial is [^1] We have re-scaled the transport coefficients ($\theta$, $\chi$, and $\lambda$) by proper factors of $T$ so that the products of these coefficients with frequency ($\omega$) or momentum ($k$) are dimensionless.
$P_{\text{BDNK}}(\omega,k)=\left[\omega(1-i\chi\,\omega)+i\lambda k^{2}\right]\times$
@@
${\mathfrak{w}}_{\rm sh}=-{i\over\theta}+i\lambda~{}k^{2}\,,\quad\omega_{\rm sh}=-i\lambda~{}k^{2}\,,$
${\mathfrak{w}}^{(1)}_{{\rm snd}}=-{i\over\theta}+{i\over 3}\left[4\lambda-{\theta\chi\over(\theta-\chi)}\right]k^{2}\,,$
${\mathfrak{w}}^{(2)}_{{\rm snd}}=-{i\over\chi}+{i\over 3}\left(\theta\chi\over\theta-\chi\right)k^{2}\,,$ (20)
$\omega^{(1)}_{{\rm snd}}={k\over\sqrt{3}}-\left(2i\lambda\over 3\right)k^{2}-\left(2\lambda^{2}\over 3\sqrt{3}\right)k^{3}\,,$
$\omega^{(1)}_{{\rm snd}}=-{k\over\sqrt{3}}-\left(2i\lambda\over 3\right)k^{2}+\left(2\lambda^{2}\over 3\sqrt{3}\right)k^{3}\,.$
@@
$R_{u}(k)=\frac{1}{k^{2}}\left({\mathfrak{w}}_{\rm sh}+{i\over\theta}\over{\mathfrak{w}}_{\rm sh}\right).$ (21)
@@
$F_{u}(\omega,k)=-{1/\theta\over{\mathfrak{w}}_{\rm sh}(\omega-{\mathfrak{w}}_{\rm sh})}.$ (22)
@@
$R_{T}(k)=i\theta\,{\mathfrak{w}}_{\rm sh}\left[F_{T}\,G_{u}\right]_{(0)}+\frac{1}{k^{2}}\left[1+{i\over\chi}\left({\mathfrak{w}}_{\rm sh}\over{\mathfrak{w}}_{{\rm snd}}^{(1)}{\mathfrak{w}}_{{\rm snd}}^{(2)}\right)\right].$ (23)
Finally, putting (23) in the same equation with $\omega$ finite, we solve for $G_{T}(\omega,k)$
@@
Note that (25) must be fulfilled only at the frequencies of the hydrodynamic sound modes, $\omega^{(1)}_{{\rm snd}}(k)$ and $\omega^{(2)}_{{\rm snd}}(k)$. Solving these equations does not determine the unknown functions $F_{T}$ and $G_{u}$ for all $\omega$ and $k$. However, our aim is to demonstrate the existence of at least one frame transformation that eliminates the non-hydrodynamic modes. By assuming $F_{T}$ and $G_{u}$ do not depend on $\omega$, equation (25) uniquely solves these functions, providing a suitable frame transformation. With this assumption, the solutions for $R_{T}$ and $G_{T}$ simplify
@@
Even after assuming $F_{T}$ and $G_{u}$ are independent of $\omega$, eq. (25) remains a set of coupled nonlinear algebraic equations, with nonlinearity due to $R_{T}$ depending on the product of $F_{T}$ and $G_{u}$. While solving is straightforward, the exact expressions for $F_{T}$ and $G_{u}$ are cumbersome compared to other functions $F_{u}$, $R_{u}$, $R_{T}$, and $G_{T}$. The complete details are presented in[29]. The final leading solutions, up to $\mathcal{O}(k^{2})$, are as follows:
@@
At this point, we note that although we provided formal exact solutions for $F_{u}$, $R_{u}$, $R_{T}$, and $G_{T}$, these should be understood as expansions in non-negative powers of $\omega$ and $k$. Polynomials of $k$ and $\omega$ in the denominator imply an infinite power series with a convergence radius computed by the zeros of these polynomials. Our study relies on the derivative expansion, and replacing $\{\omega,k^{2}\}$ with derivatives doesn’t make sense if $\omega$ and $k$ appear in the denominator. All functions $F_{u(T)}$, $G_{u(T)}$, and $R_{u(T)}$ remain finite as $\omega\to 0$ and/or $k\to 0$. This ensures no negative powers of $(u\cdot\partial)$ or $(P_{\mu}^{\nu}\partial_{\nu})$ arise after expansion and replacement. Furthermore, the final frame transformation solution must depend only on even powers of $k$ to apply the replacement without using non-analytic expressions like the square root of a derivative. Since we’re concerned with the spectrum of linearized perturbations, the order of derivatives $(u\cdot\partial)$ and $P^{\mu}_{\nu}\partial_{\mu}$ is irrelevant. The frame transformation incorporates terms of all orders in the derivative expansion, thus, the final transformed stress-energy tensor also have terms of all orders in the derivative expansion.